OpenStudy (anonymous):

I had to solve for the variable in: 3(4x-1)-2=17x+10. I got x=-1. How do I check this?

OpenStudy (ajprincess):

Plug in -1 for x in the equation and see if LHS=RHS

OpenStudy (ajprincess):

Does that help @Claudilovely?

OpenStudy (anonymous):

And that's what I did but it didn't turn out right. . . can you explain? I've been trying to figure this out for the longest.

OpenStudy (anonymous):

I do: 3(4(-1)-2=17(-1)+10 3(-5)-2=17(-1)+10 -15-2=17(-1)+10 -15-2=7 17 doesn't epual 7. What did I do wrong?

OpenStudy (ajprincess):

3(4x-1)-2=17x+10 3*4x-3*1-2=17x+10 12x-3-2=17x+10 12x-5=17x+10 Try solving for x nw nd see vat u get

OpenStudy (anonymous):

Where you get the 3 in: 3*4x-3*1-2=17x+10 ^ That one

OpenStudy (ajprincess):

3(4x-1). this means u have to multiply the terms within the parentheses by 3. That's what distributive property states. a(b-c)=a*b-a*c Getting it?

OpenStudy (anonymous):

Oh.... yup!

OpenStudy (ajprincess):

if u dnt get it plz ask:)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

wait. . .but when I got -1 for x, that was right?

OpenStudy (ajprincess):

nope. -1 is not the solution. When u plugged it and checke u didnt get LHS=RHS.

OpenStudy (anonymous):

nope, i didn't. Then can u explain what x is? So that way, maybe the checking part won't be so annoyingly hard

OpenStudy (ajprincess):

3(4x-1)-2=17x+10 3*4x-3*1-2=17x+10 12x-3-2=17x+10 12x-5=17x+10 -5=17x-12x+10 -5=5x+10 -5-10=5x -15=5x x=-15/5 x=-3 nw plug in -3 and check if LHS=RHS

OpenStudy (anonymous):

oh my! That's what I got on my first try, but I thought I was wrong. Wow. . . .

OpenStudy (anonymous):

Thank you!!!!!!!!

OpenStudy (ajprincess):

Welcome:)